OpenStudy (anonymous):

find the zeros/ complete the square 2x^2 + x - 5

OpenStudy (anonymous):

\[2x^2+x=5\] \[x^2+\frac{1}{x}x=\frac{5}{2}\] \[(x+\frac{1}{4})^2=\frac{5}{2}+\frac{1}{16}=\frac{41}{16}\] \[x+\frac{1}{4}=\pm\sqrt{41}{4}\] \[x=\frac{-1\pm\sqrt{41}}{4}\]